Suppose you sell nine July 2014 silver futures contracts on this day, at the last price of the day. Use Table 23.1 What will your profit or loss be if silver prices turn out to be $19.21 per ounce at expiration? (Do not round intermediate calculations. Enter your answer as a positive value rounded to the nearest whole number, e.g., 32.) (Click to select) What will your profit or loss be if silver prices are $19.11 per ounce at expiration? (Do not round intermediate calculations. Enter your answer as a positive value rounded to the nearest whole number, e.g., 32.) (Click to select)
Answer #1

The solution to the given question is provided below:

1.
=5000*(-9*(19.21-19.174))=-1620
Loss of $1,620

2.
=5000*(-9*(19.11-19.174))=2880
Profit of $2,880